OpenStudy (anonymous):

The solution of |3x + 12| = 18 is -10 or 2. TRUE/FALSE

OpenStudy (anonymous):

do you know what the first step to solve this is? just asking, if the answer is "no" that is ok, i will write it

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

then again you could plug in the two numbers and see if they both work that way you don't have to actually solve anything lets solve it first

OpenStudy (anonymous):

the first step is to remove the absolute value signs, and write two separate equations \[|3x+12|=18\] becomes \[3x+12=18\] or \[3x+12=-18\]

OpenStudy (anonymous):

you have to solve both separately but the steps for both are the same a) subtract \(12\) from both sides, then b) divide both sides by \(3\)

OpenStudy (anonymous):

you good with that, or still need help?

OpenStudy (anonymous):

Thank you! Im good :)

OpenStudy (anonymous):

ok good btw it is true i think